determine the kinetic energy of a 125kg roller coaster that is moving with a speed of 20.0 m/s

1
Expert's answer
2022-05-13T03:24:32-0400

The kinetic energy is given by

"E_k=\frac{mv^2}{2}"

"E_k=\rm\frac{125\: kg*(20.0\: m/s)^2}{2}=25\:000\: J"